OpenStudy (anonymous):

An engineering student stands at the edge of a cliff and throws a stone horizontally over the edge with a speed of 18 m/s. The cliff is 50 m above a flat, horizontal beach. How long after being released does the stone strike the beach below the cliff? what is its range?

OpenStudy (lgbasallote):

why couldnt it have been the mathematician :( the engineers have led a good life :'(

OpenStudy (anonymous):

it's the author's fault

OpenStudy (lgbasallote):

okay anyway...this is motion in two dimensions, yes?

OpenStudy (shane_b):

(I bet @lgbasallote goes for a quadratic here - let's see)

OpenStudy (lgbasallote):

not actually...i havent learned this stuff yet

OpenStudy (lgbasallote):

have you forgotten im in grade school? o.O

OpenStudy (lgbasallote):

i do it by splitting it into two dimensions

OpenStudy (lgbasallote):

im take the angle with least resistance

OpenStudy (lgbasallote):

i take*

OpenStudy (anonymous):

use equations of motion to solve it....

OpenStudy (shane_b):

This one is pretty straightforward. Start by solving for the time it takes to fall from a height of 50m:\[h=\frac{1}{2}{gt^2}\]Once you know that, multiply that times the horizontal velocity of 18m/s to get the range.

OpenStudy (anonymous):

sorry...busy answering some other questions....:|

OpenStudy (dls):

Use ==>S = Vit + ½ at ² (- 50) = 0 +½ (-9.8)t ² (- 50) = - ½ (9.8)t ² (4.9)t ² = 50 t ² = 50 / 4.9 .......______ t = ²√ 10.20 t = 3.19 secs

OpenStudy (dls):

Resultant velocity is given by: V = ²√ Vx² + Vy²

OpenStudy (dls):

The Vy here will be found even without using the t that was solved earlier, by using ==>2aS = Vf ² - Vi ²

OpenStudy (dls):

2(-9.8)(-50) = Vf ² - 0 Vf = 31.30 m/s = Vy ..........________ V = ²√ Vx² + Vy² ..........______________ V = ²√ (18.0)² + (31.30)² ..........________ V = ²√ 1303.69 V = 36.107 m/sec

OpenStudy (dls):

tanӨ = Vy / Vx .........=(31.30) / (18) .........=(1.74) Ө = tan‾ ¹ (1.74) Ө = 60.11°

OpenStudy (dls):

Find the range now?

OpenStudy (dls):

What is the answer?

OpenStudy (shane_b):

@DLS: You got the time right but I don't see how you're getting all that for the distance from a strictly horizontal throw.

OpenStudy (dls):

All these values to find the range.

OpenStudy (shane_b):

The horizontal velocity is a constant 18 m/s...until it hits the ground.

OpenStudy (shane_b):

\[(18m/s)(3.19s)=57.42m\]

OpenStudy (dls):

noooo,it is thrown at some angle?

OpenStudy (shane_b):

I don't see how there's an angle involved since it says "throws a stone horizontally over the edge "

OpenStudy (dls):